"I think we goofed when we hired that new assistant controller," said Ruth Scarpino, president of Provost Industries. "Just look at this report that he prepared for last month for the Finishing Department. I can’t understand it."

Finishing Department costs:
Work in process inventory, April 1, 450 units; materials
100% complete; conversion 60% complete
$ 8,208 *
Costs transferred in during the month from the
preceding department, 1,950 units
17,940
Materials cost added during the month 6,210
Conversion costs incurred during the month 13,920
Total departmental costs $ 46,278
Finishing Department costs assigned to:
Units completed and transferred to finished goods,
1,800 units at $25.71 per unit
$ 46,278
Work in process inventory, April 30, 600 units;
materials 0% complete; conversion 35% complete
0
Total departmental costs assigned $ 46,278

*Consists of cost transferred in, $4,068; materials cost, $1,980; and conversion cost, $2,160.

"He's struggling to learn our system," replied Frank Harrop, the operations manager. "The problem is that he's been away from process costing for a long time, and it's coming back slowly."

"It's not just the format of his report that I'm concerned about. Look at that $25.71 unit cost that he's come up with for April. Doesn't that seem high to you?" said Ms. Scarpino.

"Yes, it does seem high; but on the other hand, I know we had an increase in materials prices during April, and that may be the explanation," replied Mr. Harrop. "I’ll get someone else to redo this report and then we can see what’s going on."

Provost Industries manufactures a ceramic product that goes through two processing departments—Molding and Finishing. The company uses the weighted-average method in its process costing.

Required:

1-a. Calculate the equivalent units of production.

1-b. Calculate the cost per equivalent unit. (Round your answers to 2 decimal places.)

1-c. How much cost should have been assigned to the ending work in process inventory? (Round your intermediate calculations to 2 decimal places.)

1-d. How much cost should have been assigned to the units completed and transferred to finished goods? (Round your intermediate calculations to 2 decimal places.)
Answer #1

1-a)

Transfered in Material conversion
units completed and trasfered out 1800 1800 1800
Ending WIP 600 0 210      [600*.35]
Equivalent units (ANs 1a) 2400 1800 2010
Total cost incurred till date 4068+17940=22008 1980+6210=8190 2160+13920=16080
cost per equivalent unit (Ans 1 b) 22008/2400=$ 9.17 8190/1800=$ 4.55 16080/2010=8

1-c) cost should have been assigned to the ending work in process inventory :[600*9.17]+[0*4.55]+[210*8]

           = 5502+ 0+ 1680

            = $ 7182

1d) cost should have been assigned to the units completed and transferred to finished goods:[1800*9.17]+[1800*4.55]+[1800*8]

= 16506+ 8190+ 14400

= $ 39096